2 times the sq root of 90 + 3 times the sq root of 40 - 4 times the sq root of 10 = ??
Answer by lazi(15)

x^(1/2) = square root of x
2(90)^(1/2) + 3(40)^(1/2) - 4(10)^(1/2) =
to add or subtract them we need to make sure they have the same square root; its like having the same terms. (i.e. 2(x) + 3(x) = 5(x)...)
like finding the common denominator we look for what number they have in common.
90 = 9 * 10
40 = 4 * 10
10 = 10
9 = 3^2
4 = 2^2
2(3^2 * 10)^(1/2) + 3(2^2 *10)^(1/2) - 4(10)^(1/2) =
[2 * 3 * (10)^(1/2)] + [3 * 2 * (10)^(1/2)] - [4 * (10)^(1/2)] =
6(10)^(1/2) + 6(10)^(1/2) - 4(10)^(1/2) =
now it's just simply adding and subtracting the number before (10)^(1/2)
6+6-4= 12-4=8
the answer is 8(10)^(1/2) ; 8 times the square root of 10